OpenStudy (anonymous):

Use the function rule ƒ(x) = x^2 – 5x + 1. Find the output ƒ(–3). A.7 B.–5 C.25 D.5 Please help me, i've been stuck on this for a while and have no clue how to do the first step. Can someone give me explaination?

OpenStudy (anonymous):

Sure... See f(-3) means simply that you will put x = -3 there in your expression \(x^2 - 5x + 1\)..

OpenStudy (stamp):

f(x) : f(-3) ---> x = -3

OpenStudy (anonymous):

So, you will get: \[f(x) = x^2 -5x + 1\] Put x = -3 here: \[f(-3) = (-3)^2 - 5(-3) + 1\] Solve now on your own..

OpenStudy (stamp):

f(x) = x^2 -5x + 1, f(-3) ---> x = -3 f(-3) = (-3)^2 - 5(-3) + 1

OpenStudy (anonymous):

C. 25?

OpenStudy (stamp):

Whats with the question mark??

OpenStudy (anonymous):

Asking if that the answer.

OpenStudy (stamp):

(-3)^2 -5(-3) + 1 = 9 + 15 + 1 = 25 period.

OpenStudy (anonymous):

Have faith in yourself..

OpenStudy (anonymous):

Yes, you are right...

OpenStudy (anonymous):

Oh ok this makes sense, thank you very much guys.

OpenStudy (stamp):

@kard192 np

OpenStudy (anonymous):

You are welcome dear..